OpenStudy (anonymous):

subtract (10x^2+7x-2) - (4x^2-2x+2)

OpenStudy (anonymous):

6x^2+9x-4

OpenStudy (akshay_budhkar):

6x^2+9x-4 just subtract the like terms

OpenStudy (anonymous):

first remove the parentheses distributing the minus sign. change the sign of everything after the parentheses in the second term. you get \[10x^2+7x-2-4x^2+2x-2\]

OpenStudy (anonymous):

yup. gone now!

OpenStudy (akshay_budhkar):

lol yea

OpenStudy (anonymous):

rosie you see how to distribute the minus sign. you turn all + to - and all - to +

OpenStudy (anonymous):

yes i understand that

OpenStudy (anonymous):

\[- (4x^2-2x+2)=-4x^2+2x-2\] leave the first term alone

OpenStudy (anonymous):

so once you have \[10x^2+7x-2-4x^2+2x-2\] you need to combine like terms. can you do that?

OpenStudy (anonymous):

yes

OpenStudy (akshay_budhkar):

i like the way you explain step by step...

OpenStudy (anonymous):

i will group them together. they are \[10x^2-4x^2+7x+2x-2-2\] go ahead and combine them

OpenStudy (anonymous):

@akshay thanks

OpenStudy (anonymous):

6x^2 +9x

OpenStudy (anonymous):

close

OpenStudy (akshay_budhkar):

you are close but check again

OpenStudy (anonymous):

\[10x^2-4x^2=6x^2\] \[7x+2x=9x\] \[-2-2\]=?

OpenStudy (anonymous):

you made a common mistake on the last one. \[-2-2\] is not the same as \[2-2\]

OpenStudy (anonymous):

\[2-2=0\] but \[-2-2=?\]

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

got it so now you have the answer!

OpenStudy (akshay_budhkar):

cool you got it!!! keep it up!

OpenStudy (anonymous):

careful with your signs on the test. don't rush through. really.

OpenStudy (anonymous):

Multiply (4x+5)(2x^2-x-3

OpenStudy (anonymous):

think we did this one yes? multiply everything in the second parentheses by 4x, then multiply it all by 5, then combine like terms

OpenStudy (anonymous):

whats the rule for the signs? multiply and division are the same and addition and subtraction are the same?

OpenStudy (anonymous):

i am not sure what you mean. the rules for multiplication and division are different from addition and subtraction

OpenStudy (anonymous):

in multiplication and addition two minuses make a plus, but not so in addition. add to negative numbers, get a negative number.

OpenStudy (anonymous):

that doesnt make any sense???

OpenStudy (anonymous):

(1/4 a-b)(1/4a+b)

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

sure it does. think about it like this. if you spend $5 and then spend $10 that is like \[-5-10\] which is clearly \[-15\] not 15

OpenStudy (anonymous):

if you lose money twice it is not like you gain money